Apply Newtons Method to approximate the x-value(s) of the indicated point(s) of intersection of the two graphs. Continue theApply Newton's Method to approximate the x-value(s) of the indicated point(s) of intersection of the two graphs. Continue the iterations until two successive approximations differ by less than 0.001. [Hint: Let

h(x) = f(x) − g(x).]

f(x) = 2x + 2

g(x) =

x + 10

find x please
Answer #1


liven, 22 +2 F(x) = g(x) 5x + 10 h(x) = f (x) - g(x) 2 x + 2 - {x+10 We know that, for h(x)=0 formula for Mewtons method is x2 = 0.630 2 x 0.631 +2 -0.631 +10 2 2 Jo.661+10 - 0.630 Xz = x2 f (x2) J(x2) 2*0-630+2-10.630110 0.630 2- bast 250.630 +10